The measures of the interior angles of a convex octagon are 45x, 40x, 155, 120, 155, 38x, 158, and 41x. What is the measure of the smallest interior angle?
Answer by nycsharkman(136)

The sum of the interior angles of an octagon is 1080 degrees.
You were given:
45x, 40x, 155, 120, 155, 38x, 158, and 41x. You were told to find the smallest angle. The smallest angle is 38x but we need to find x first.
We add what has been given, equate to 1080 and solve for x.
45x + 40x + 155 + 120 + 155 + 38x + 158 + 41x = 1080

164x + 588 = 1080
Subtract 588 from 1080.
164x = 1080 - 588
164x = 492
Divide both sides by 164.
x = 492/164
x = 3
We now multiply 3 times 38 to find the smallest angle.
Smallest angle = 3 times 38
Smallest angle = 114
Is this clear?